you invested $18,000 in two accounts paying 5% and 9% annual interest. if the total interest earned for the year was $1380, how much was invested at each rate?

Answer by richwmiller(17219)

x+y=18000,
0.05*x+0.09*y=1380
x=18000-y
0.05*(18000-y)+0.09*y=1380
900.0-0.05y+0.09*y=1380
0.04*y=480.0
y=12000.0
x=18000-y
x=6000.0 y=12000.0